X Y Tot
A 24 31 55
B 13 66 79
Tot 37 97 134
Is P(X or A) = 68/134?
and
Is P(X|A) = 24/55?

Answer by stanbon(75887)

X Y Tot
A 24 31 55
B 13 66 79
Tot 37 97 134
Is P(X or A) = (37 + 55 - 24)/134 = 68/134?
and
Is P(X|A) = P(X and A)/P(A) = 24/55?
----
Both are correct.
Cheers,
Stan H.
---------------